# Token-Importance Guided Direct Preference Optimization

### v1

@@ Appendix A Theoretical Proof
A.1 Proof of Lemma 4.1
Using the linear property of expectation, the expectation of the weighted reward difference between the preferred response $y_{\text{w}}$ and the non-preferred response $y_{\text{l}}$ can be decomposed as:
$\mathbb{E}[\Delta r_{\text{token}}]$ $=\mathbb{E}\left[\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})-\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})\right]$ (18)
$=\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]-\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})].$
where $T_{w}$ and $T_{l}$ are the token lengths of the preferred and non-preferred responses respectively, and $w_{t}^{w}$ and $w_{t}^{l}$ are the token-importance weights calculated dynamically through gradient attribution. For the token $t$ in the preferred response, since the weight $w_{t}^{w}$ is positively correlated with the reward $r_{\phi}(x,y_{w}^{t})$, i.e., $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\geq 0$, then we have
$\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]=\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})]+\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))$ (19)
Similarly, for $y_{\text{l}}$, the weight $w^{\text{l}}_{t}$ is negatively correlated with its reward $r_{\phi}(x,y_{\text{l}}^{t})$, i.e., $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\leq 0$. Then, we have:
$\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})]=\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})]+\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))$ (20)
Substitute the positive correlation condition $\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))\geq 0$ into (19), we obtain
$\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}r_{\phi}(x,y_{\text{w}}^{t})]\geq\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})],$ (21)
Similarly, with $\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))\leq 0$ for (20), we have:
$\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}r_{\phi}(x,y_{\text{l}}^{t})]\leq\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})].$ (22)
Denote $\lambda=\min\big{\{}\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t},\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}\big{\}}>1$, we have
$\sum_{t=1}^{T_{\text{w}}}\mathbb{E}[w^{\text{w}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{w}}^{t})]\geq\lambda\mathbb{E}\left[\sum_{t=1}^{T_{\text{w}}}r_{\phi}(x,y_{\text{w}}^{t})\right]=\lambda\mathbb{E}[r_{\phi}(x,y_{\text{w}})],$ (23)
and
$\sum_{t=1}^{T_{\text{l}}}\mathbb{E}[w^{\text{l}}_{t}]\mathbb{E}[r_{\phi}(x,y_{\text{l}}^{t})]\geq\lambda\mathbb{E}\left[\sum_{t=1}^{T_{\text{l}}}r_{\phi}(x,y_{\text{l}}^{t})\right]=\lambda\mathbb{E}[r_{\phi}(x,y_{\text{l}})].$ (24)
Substituting (21) and (22) into (18), there is
$\mathbb{E}[\Delta r_{\text{token}}]$ $\geq\lambda\mathbb{E}[r_{\phi}(x,y_{\text{w}})]+\sum_{t=1}^{T_{\text{w}}}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))-\left(\lambda\mathbb{E}[r_{\phi}(x,y_{\text{l}})]-\sum_{t=1}^{T_{\text{l}}}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|\right)$ (25)
$=\lambda\mathbb{E}[\Delta r_{\text{global}}]+\underbrace{\sum_{t=1}^{T_{\text{w}}}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))+\sum_{t=1}^{T_{\text{l}}}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|}_{C>0}.$
Let $C=\sum_{t=1}^{T}\text{Cov}(w^{\text{w}}_{t},r_{\phi}(x,y_{\text{w}}^{t}))+\sum_{t=1}^{T}|\text{Cov}(w^{\text{l}}_{t},r_{\phi}(x,y_{\text{l}}^{t}))|$, $C>0$ and define $\alpha=\lambda+\frac{C}{\mathbb{E}[\Delta r_{\text{global}}]}.$ Since the global reward of the preferred response $y_{\text{w}}$ is necessarily higher than that of the non-preferred response $y_{\text{l}}$, i.e., $\mathbb{E}[\Delta r_{\text{global}}]>0$, we have $\alpha>1$, and
$\mathbb{E}[\Delta r_{\text{token}}]\geq\alpha\cdot\mathbb{E}[\Delta r_{\text{global}}].$ (26)
Hence, we show the Lemma 4.1.
A.2 Proof of Theorem 4.2
Since the sigmoid function $\sigma(x)=1/(1+\exp(-x))$ is convex, using Jensen’s inequality to $\mathcal{L}_{\text{DPO-w}}$, we have
$\mathcal{L}_{\text{DPO-w}}=-\mathbb{E}\left[\log\sigma\left(\beta\Delta r_{\text{token}}\right)\right]\leq-\log\sigma\left(\beta\mathbb{E}\left[\Delta r_{\text{token}}\right]\right)$ (27)
According to the assumption $\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right]$ shown in Appendix 1, with the monotonicity of $\sigma(z)$, we have
$-\log\sigma\left(\beta\mathbb{E}\left[\Delta r_{\text{token}}\right]\right)\leq-\log\sigma\left(\beta\alpha\mathbb{E}\left[\Delta r_{\text{global}}\right]\right).$ (28)
Denote $z=\beta\mathbb{E}\left[\Delta r_{\text{global}}\right]>0$ and construct the function $g(\alpha)=\log(1+e^{-\alpha z})+(\alpha-1)z$. Since the monotonicity and $g(1)=\log(1+e^{-z})$, when $\alpha>1$, we have $g(\alpha)<g(1)$, i.e.,
$-\log\sigma(\alpha z)=\log(1+e^{-\alpha z})\leq\log(1+e^{-z})-(\alpha-1)z=-\log\sigma(z)-(\alpha-1)z$ (29)
Substituting $z=\beta\mathbb{E}\left[\Delta r_{\text{global}}\right]$ into (29), we have
$\mathcal{L}_{\text{DPO-w}}\leq\mathcal{L}_{\text{DPO}}-\beta(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ (30)
According to Eq.(14), when the optimization is close to being completed,
$\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}$ (31)
$-\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}\geq-\alpha$
and there exists a constant $\Delta_{\text{triplet}}$ such that $\mathcal{L}_{\text{triplet}}\leq\Delta_{\text{triplet}}$. Assuming $\gamma<\beta$ and combining the importance-weighted DPO loss and triplet loss, we have
$\mathcal{L}_{\text{TI-DPO}}$ $=\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}}$ (32)
$\leq\mathcal{L}_{\text{DPO}}-\beta(\alpha-1)\mathbb{E}[\Delta r_{\text{global}}]+\beta\Delta_{\text{triplet}}$
Let $\Delta_{\text{triplet}}=\frac{1}{2}(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ and substitute into (32), then we have
$\mathcal{L}_{\text{TI-DPO}}$ $\leq\mathcal{L}_{\text{DPO}}-\frac{1}{2}\beta(\alpha-1)\mathbb{E}\left[\Delta r_{\text{global}}\right]$ (33)
$\leq\mathcal{L}_{\text{DPO}}-\beta\Delta_{\text{triplet}}$
Hence, Theorem 4.2 holds.
A.3 Gradient Analysis of Loss
@@
$\nabla_{\theta}\Delta r_{\text{token}}=\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$ (35)
@@
$\nabla_{\theta}\mathcal{L}_{\text{DPO-w}}=-\beta(1-\sigma(\beta\Delta r_{\text{token}}))\left[\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})\right].$ (36)
@@
$\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}=$ $-\beta(1-\sigma(\beta\Delta r_{\text{token}}))\left[\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})\right]$ (42)
$+2\gamma\left[\sum_{t=1}^{T_{\text{w}}}(d_{t}-b_{t})(\nabla_{\theta}d_{t}-\nabla_{\theta}b_{t})-\sum_{t=1}^{T_{\text{l}}}(c_{t}-d_{t})(\nabla_{\theta}c_{t}-\nabla_{\theta}d_{t})\right]$
@@ B.1 Distribution of Weights
Figure 4 visualizes the distribution of importance weights assigned to tokens in different tasks by TI-DPO with histograms and box plots, intuitively explaining how the importance mechanism dynamically focuses on key tokens according to task characteristics. The y-axis of the histogram is token frequency, which represents the number of occurrences of tokens with varying importance. The box plots in the upper right corner are used to visualize the distribution of token frequencies for different weights, where the red line represents the median, reflecting the concentration of token frequencies. In the GSM8K and GPQA datasets, with only a few key symbols crucial for answering, the importance weights mostly concentrate in the interval of $[0.2,0.5]$ approximately. For the TruthfulQA and IFEval datasets, token distribution is concentrated in the $[0.6,0.8]$ weight interval. The MMLU and HumanEval datasets cover a wide range of content, and the model assigns diverse weights to various tokens.
@@
Figure 4: Distribution patterns of gradient-based token importance weights in six benchmark tasks (GSM8K, TruthfulQA, MMLU, GPQA, HumanEval, IFEval)
B.2 Case Demo
In the TI-DPO framework, response A corresponds to the preferred response $y_{\text{w}}$ in the dataset, which represents the high-quality, human-preferred output aligned with safety, professionalism, and correctness (e.g., "seek medical attention promptly" in the medical case). These tokens are assigned high importance weights to prioritize critical elements in human judgments. As show in Table 5, critical tokens like "seek" (0.93), "medical" (0.87), "attention" (0.85), "promptly" (1.00), "avoid" (0.92), and "self-medicating" (0.89) receive high importance weights, reflecting their role in ensuring safety and compliance with medical standards. These weights act as a "spotlight" to prioritize tokens that most influence human judgments, such as emergency actions and avoidance of self-treatment.
Response C in Table 5 corresponds to the less preferred response $y_{\text{l}}$, representing low-quality or risky outputs that deviate from human preferences (e.g., "take painkillers casually" in the example). The model includes high-risk tokens like "painkillers" (0.91), "casually" (1.00), and "fine" (0.97), which receive peak weights due to their potential to mislead users into unsafe self-medication. TI-DPO’s gradient-based attribution mechanism identifies these tokens as critical for preference misalignment, suppressing their influence during generation.
In Table 5, response B represents an intermediate generated response (e.g., "get more rest… consult a doctor"), which is neither the top preferred nor the worst case. In the triplet loss structure, B acts as an anchor that is guided to approach $y_{\text{w}}$ (A) and distance from $y_{\text{l}}$ (C). Key tokens like "worsen" (0.88), "should" (0.82), "consult" (0.90), and "doctor" (0.95) have elevated weights but are less intense than those in A, indicating their secondary importance in guiding less urgent but still reasonable advice. By incorporating B, TI-DPO promotes more nuanced optimization, where intermediate outputs are refined to better match human preferences through token-level importance weights and triplet constraints.
Table 3: Token Importance
Assignment of A
Token Weight
Based 0.05
on 0.05
your 0.05
symptoms 0.18
it 0.03
is 0.03
recommended 0.20
that 0.02
you 0.02
seek 0.93
medical 0.87
attention 0.85
promptly 1.00
and 0.03
avoid 0.92
self-medicating 0.89
Table 4: Token Importance
Assignment of B
Token Weight
According 0.04
to 0.04
your 0.04
description 0.07
it 0.03
is 0.03
advised 0.13
to 0.02
get 0.06
more 0.06
rest 0.11
symptoms 0.18
worsen 0.88
you 0.03
should 0.82
consult 0.90
doctor 0.95
Table 5: Token Importance
Assignment of C
Token Weight
Don’t 0.21
worry 0.18
you 0.04
can 0.04
just 0.09
take 0.09
some 0.05
painkillers 0.91
casually 1.00
it 0.02
should 0.06
be 0.03
fine 0.97
B.3 Other Tables

### v2

@@ Appendix A Theoretical Proof
A.1 Proof of Lemma 1
In standard DPO, the optimization objective implicitly assigns a uniform weight of $1$ to all tokens. Consequently, the estimator aggregates the noise from the entire non-critical set. Assuming the noise terms $\epsilon_{t}$ are independent, the total variance for DPO is the sum of their individual variances:
$\sigma^{2}_{\text{DPO}}=\sum_{t\in\mathcal{N}}1^{2}\cdot\text{Var}(\epsilon_{t})=|\mathcal{N}|\sigma^{2}_{\epsilon},$ (17)
where $|\mathcal{N}|$ denotes the number of non-critical tokens.
In TI-DPO, the contribution of each token is scaled by its importance weight $w_{t}$. The variance of the weighted estimator becomes:
$\sigma^{2}_{\text{TI-DPO}}=\sum_{t\in\mathcal{N}}w_{t}^{2}\cdot\text{Var}(\epsilon_{t})=\sigma^{2}_{\epsilon}\sum_{t\in\mathcal{N}}w_{t}^{2}$ (18)
Given the condition that the weights in the non-critical region are suppressed ($w_{t}^{2}<1$), the sum of the squared weights is strictly less than the count of the tokens:
$\sum_{t\in\mathcal{N}}w_{t}^{2}<\sum_{t\in\mathcal{N}}1=|\mathcal{N}|$ (19)
Substituting the definitions of $\sigma^{2}_{\text{TI}}$ and $\sigma^{2}_{\text{DPO}}$, we conclude:
$\sigma^{2}_{\text{TI-DPO}}<\sigma^{2}_{\text{DPO}}$ (20)
A.2 Proof of Theorem 2
As shown in Eq.(3), the loss function for DPO becomes: $\mathcal{L}_{\text{DPO}}=\mathbb{E}[-\log\sigma(\beta\Delta r_{\text{global}})].$ Similarly, the TI-DPO loss is $\mathcal{L}_{\text{TI-DPO}}=\mathbb{E}[-\log\sigma(\beta\Delta r_{\text{token}})]$. Thus, we define the generic scalar loss function for an estimator $R$ (where $R\in\{\Delta r_{\text{global}},\Delta r_{\text{token}}\}$) as $f(R)=-\log\sigma(\beta R)$.
We analyze the properties of $f(R)$. The first and second derivatives with respect to $R$ are:
$f^{\prime}(R)=-\beta(1-\sigma(\beta R))$ (21)
$f^{\prime\prime}(R)=\beta^{2}\sigma(\beta R)(1-\sigma(\beta R))$ (22)
Since the sigmoid function satisfies $\sigma(\cdot)\in(0,1)$, the second derivative $f^{\prime\prime}(R)$ is strictly positive for any finite $R$. Therefore, $f(R)$ is a strictly convex function. Let $\kappa=\inf f^{\prime\prime}(\xi)>0$ denote the lower bound of the curvature within the effective optimization region.
We employ a second-order Taylor expansion of the loss function $f(R)$ around the true (expected) reward difference $\mu$. Let $R$ represent the estimator (either $\Delta r_{\text{global}}$ or $\Delta r_{\text{token}}$). The expansion with the Lagrange remainder is:
$f(R)=f(\mu)+f^{\prime}(\mu)(R-\mu)+\frac{1}{2}f^{\prime\prime}(\xi)(R-\mu)^{2},$ (23)
where $\xi$ lies between $R$ and $\mu$. Taking the expectation $\mathbb{E}[\cdot]$ on both sides, and applying the unbiasedness assumption $\mathbb{E}[R]=\mu$ (which eliminates the linear term):
$\mathbb{E}[f(R)]=f(\mu)+\frac{1}{2}\mathbb{E}[f^{\prime\prime}(\xi)(R-\mu)^{2}].$ (24)
Using the curvature lower bound $\kappa$, we can lower-bound the expected loss in terms of the estimator’s variance $\sigma^{2}=\text{Var}(R)=\mathbb{E}[(R-\mu)^{2}]$:
$\mathbb{E}[f(R)]\geq f(\mu)+\frac{1}{2}\kappa\sigma^{2}.$ (25)
The difference is approximated by the difference in variances:
$\mathcal{L}_{\text{DPO}}-\mathcal{L}_{\text{TI-DPO}}\approx\frac{1}{2}\mathbb{E}[f^{\prime\prime}(\xi)](\sigma^{2}_{\text{DPO}}-\sigma^{2}_{\text{TI}})$ (26)
Applying the lower bound $\kappa$ and substituting the strictly positive variance reduction term $\Delta_{\sigma^{2}}=\sigma^{2}_{\text{DPO}}-\sigma^{2}_{\text{TI-DPO}}$ derived in Lemma 1, we obtain:
$\mathcal{L}_{\text{DPO}}-\mathcal{L}_{\text{TI-DPO}}\geq\frac{1}{2}\kappa\Delta_{\sigma^{2}}.$ (27)
Since $\kappa>0$ and $\Delta_{\sigma^{2}}>0$, this completes the proof.
A.3 Proof of Theorem 3
Proof.
We begin by defining the Sparse Criticality Assumption: For a given input $x$, the token indices of a response $y$ partition into a critical set $\mathcal{C}$ and a non-critical set $\mathcal{N}$, where $|\mathcal{C}|\ll|\mathcal{N}|$. The true reward function $r^{*}(x,y)$ depends solely on tokens in $\mathcal{C}$. Consequently, any deviation from the reference model $\pi_{\text{ref}}$ on tokens in $\mathcal{N}$ incurs a KL cost without yielding any reward gain.
The total KL divergence constraint $K_{\text{total}}$ decomposes token-wise via the chain rule:
$D_{KL}(\pi||\pi_{\text{ref}})=\sum_{t=1}^{T}\mathbb{E}_{y^{<t}\sim\pi}\left[D_{KL}(\pi(\cdot|y^{<t},x)||\pi_{\text{ref}}(\cdot|y^{<t},x))\right]=K_{\mathcal{C}}+K_{\mathcal{N}},$ (28)
where $K_{\mathcal{C}}$ and $K_{\mathcal{N}}$ represent the KL divergence allocated to critical and non-critical tokens, respectively.
In standard DPO, the optimal policy takes the form $\pi_{\text{DPO}}(y|x)\propto\pi_{\text{ref}}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)$. Since the implicit reward model $r_{\phi}$ is optimized at the sequence level, it inevitably captures spurious correlations, distributing non-zero gradients across all tokens, including those in $\mathcal{N}$. This results in “KL divergence waste”, where $\pi_{\text{DPO}}$ diverges from $\pi_{\text{ref}}$ on non-critical tokens, implying $K_{\mathcal{N}}^{\text{DPO}}\geq\epsilon$ for some $\epsilon>0$. The effective KL divergence available for critical tokens is thus limited to $K_{\mathcal{C}}^{\text{DPO}}\leq K_{\text{total}}-\epsilon$.
In contrast, TI-DPO incorporates token importance weights $w_{t}$. We assume the weights align with the sparsity structure, such that $w_{t}\approx 0$ for $t\in\mathcal{N}$ (due to low gradient attribution from non-critical tokens). This weighting suppresses the update signal on non-critical tokens, causing the policy to default to the reference, i.e., $\pi_{\text{TI-DPO}}(\cdot|y^{<t},x)\approx\pi_{\text{ref}}(\cdot|y^{<t},x)$ for $t\in\mathcal{N}$. Consequently, TI-DPO minimizes the wasted KL divergence, $K_{\mathcal{N}}^{\text{TI-DPO}}\approx 0$, allowing nearly the full constraint to be allocated to the critical set: $K_{\mathcal{C}}^{\text{TI-DPO}}\approx K_{\text{total}}$.
Since the maximum achievable expected reward $f(K)$ is a strictly increasing and concave function of the KL divergence allocated to reward-relevant tokens (a property derived from the rate-distortion nature of the RL objective), the larger effective allocation of TI-DPO implies a higher upper bound on the expected reward. Specifically,
$\mathbb{E}_{\pi_{\text{TI-DPO}}}[r^{*}]-\mathbb{E}_{\pi_{\text{DPO}}}[r^{*}]=f(K_{\mathcal{C}}^{\text{TI-DPO}})-f(K_{\mathcal{C}}^{\text{DPO}})\geq f(K_{\text{total}})-f(K_{\text{total}}-\epsilon)\triangleq\delta>0.$ (29)
This confirms that TI-DPO achieves a higher expected reward by efficiently reallocating the KL constraint. ∎
A.4 Gradient Analysis of Loss
@@
$\nabla_{\theta}\Delta r_{\text{token}}$ $=\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}$ (31)
$-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$
@@
$\nabla_{\theta}\mathcal{L}_{\text{DPO-w}}=-\beta(1-\sigma(\beta\Delta r_{\text{token}})){[}\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})$ (32)
$-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t}){]}.$
@@
$\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}=$ $-\beta(1-\sigma(\beta\Delta r_{\text{token}})){[}\sum_{t=1}^{T_{\text{w}}}w_{t}^{\text{w}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})-\sum_{t=1}^{T_{\text{l}}}w_{t}^{\text{l}}\nabla_{\theta}\log\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t}){]}$ (38)
$+2\gamma{[}\sum_{t=1}^{T_{\text{w}}}(d_{t}-b_{t})(\nabla_{\theta}d_{t}-\nabla_{\theta}b_{t})-\sum_{t=1}^{T_{\text{l}}}(c_{t}-d_{t})(\nabla_{\theta}c_{t}-\nabla_{\theta}d_{t}){]}$
Table A1: Distribution of Token Importance Weight, Performance Improvement, and Sample-level Pearson Correlation Coefficient in Each Task
Task Q1 Q2 (Median) Q3 Average weight Performance Improvement $\Delta_{\text{ACC}}$ (%) Sample-level Pearson $r$ Top-5 weight vs accuracy rate
GSM8K 0.22 0.33 0.45 0.34 +4.7 0.29
GPQA 0.18 0.28 0.42 0.30 +5.0 0.22
TruthfulQA 0.70 0.75 0.85 0.78 +5.3 0.31
IFEval 0.68 0.75 0.85 0.77 +5.7 0.27
MMLU 0.30 0.50 0.70 0.50 +4.7 0.18
HumanEval 0.48 0.60 0.70 0.59 +6.0 0.35
@@ B.1 Distribution of Weights
Figure B1 visualizes the distribution of importance weights assigned to tokens in different tasks by TI-DPO with histograms and box plots, intuitively explaining how the importance mechanism dynamically focuses on key tokens according to task characteristics. The y-axis of the histogram is token frequency, which represents the number of occurrences of tokens with varying importance. The box plots in the upper right corner are used to visualize the distribution of token frequencies for different weights, where the red line represents the median, reflecting the concentration of token frequencies. In the GSM8K and GPQA datasets, with only a few key symbols crucial for answering, the importance weights mostly concentrate in the interval of $[0.2,0.5]$ approximately. For the TruthfulQA and IFEval datasets, token distribution is concentrated in the $[0.6,0.8]$ weight interval. The MMLU and HumanEval datasets cover a wide range of content, and the model assigns diverse weights to various tokens.
@@
Figure B1: Distribution patterns of gradient-based token importance weights in six benchmark tasks (GSM8K, TruthfulQA, MMLU, GPQA, HumanEval, IFEval
B.2 Pearson Correlation Coefficient
Table A1 presents multiple metrics for six tasks, namely GSM8K, GPQA, TruthfulQA, IFEval, MMLU, and HumanEval, including Q1, Q2 (median), Q3, average weight (mid-point of the interval), performance improvement $\Delta_{\text{ACC}}$ (%), and sample-level Pearson $r$ (Top 5 weights $vs$ accuracy). Among them, the “sample-level Pearson rr” is calculated based on the average weights of the Top-5 tokens for 100 questions in each task and the correctness of the answers to these questions (binary values), which reflects the microscopic internal correlation. The performance improvement $\Delta_{\text{ACC}}$ is calculated from the average accuracy gain of TI-DPO compared to DPO. The calculation process is divided into the calculation of microscopic and macroscopic Pearson correlation coefficients:
  • Microscopic calculation: To calculate the sample-level Pearson correlation coefficient, first, randomly sample 100 samples from the test set without replacement for each task. Subsequently, for each sample, use the TI-DPO model to perform forward and backward propagation to calculate the importance weight of each token, extract the Top-5 tokens with the highest weights, and calculate their average. Then, record the consistency between the model’s answer and the standard answer for this sample, marking it as 1 when correct and 0 when incorrect. Finally, substitute the obtained sequence of average weights of the Top-5 tokens and the sequence of answer correctness into the Pearson correlation coefficient Eq.(39) for calculation, where $\bar{w}$ and $\bar{\Delta}$ in the formula represent the means of the two sets of data, respectively.
    $r=\frac{\sum_{i}\left(w_{i}-\overline{w}\right)\left(\Delta_{i}-\overline{\Delta}\right)}{\sqrt{\sum_{i}\left(w_{i}-\overline{w}\right)^{2}}\sqrt{\sum_{i}\left(\Delta_{i}-\overline{\Delta}\right)^{2}}}$ (39)
    The results show that the correlation coefficient $r\approx 0.35$ for the HumanEval task, which is the highest among all tasks, indicating that in the code generation task, the correlation between token importance and the probability of answering correctly is the strongest. On the other hand, the correlation coefficient for the MMLU task is the lowest, approximately 0.18, suggesting that in the multi-task and multi-disciplinary test, the relationship between token importance and the correctness of a single question is relatively weak.
  • Macroscopic calculation: Based on the average weights (approximated by the median) and performance improvements of the 6 tasks, first construct the average weight vector: $w=[0.33,0.28,0.75,0.75,0.50,0.60]$ (Order: GSM8K, GPQA, TruthfulQA, IFEval, MMLU, HumanEval), and the performance improvement vector $\Delta=[4.7,5.0,5.3,5.7,4.7,6.0]$. Then substitute them into the Pearson correlation formula (Eq.(39)). In the specific calculation steps, first obtain $\bar{w}\approx 0.535$ and $\bar{\Delta}\approx 5.4$, and then substitute each value of the two sets of vectors into the formula in turn of summation and square-root operations. The final result shows that the overall Pearson correlation coefficient $r\approx 0.65$ at the task level, indicating a moderately strong positive correlation between the average token importance of the six tasks and the performance improvement.
The coefficient difference between the overall correlation at the task level and the correlation at the single-task sample level stems from their fundamental differences at the computational level. The overall correlation at the task level takes six tasks as the sample size and analyzes the corresponding relationship between the “average token weight” of each task and the “overall performance improvement”, essentially reflecting the macro correlation across tasks. Due to the significant differences in the weight distribution centers and improvement amplitudes of different tasks, this “inter-task difference” tends to magnify the correlation, resulting in a correlation coefficient $r\approx 0.65$.
In contrast, the correlation at the single-task sample level takes 100 questions in each task as the sample size, focusing on the correlation between the “average Top-5 token weight” of each question within the same task and the correctness of the answer to that question. Since the samples fall within the same distribution range, the signal between the weight and the correctness of the answer is weak, and it is affected by noises such as the diversity of prompts, fluctuations in question difficulty, and the randomness of token gradients. Therefore, the correlation coefficient is only between 0.18 and 0.35. This result is reasonable: the overall correlation at the task level indicates that TI-DPO has a more significant improvement on tasks with concentrated weight distributions. The weak micro-correlation at the single-task level shows that token importance is only one of the factors affecting the correctness of a single question.
B.3 Experimental Results Across Base Models